OpenStudy (anonymous):

Prove the identity: 2cos^(2)(x/2)-1=cosx

OpenStudy (raden):

use half angle formula : cos^(2)(x/2) = (1 + cos(2 * x/2))/2 = (1+cosx)/2

OpenStudy (raden):

got it ?